# Clarification and Coinduction of Tambara Functors

### v1

@@ Abstract.
Tambara functors are equivariant analogues of rings arising in representation theory and equivariant homotopy theory. We introduce the notion of a clarified Tambara functor and show that under mild conditions every Tambara functor admits a decomposition as a product of coinductions of clarified Tambara functors; projection onto the non-coinduced part defines a reflective localization we call clarification. Through this perspective we study Nakaoka spectra, Morita invariance and $K$-theory of Tambara functors, field-like and domain-like Tambara functors, $G$-$\mathbb{E}_{\infty}$-rings, Nullstellensatzian clarified Tambara functors, and transversal Tambara functors.
@@ 1. Introduction
We focus our attention on $G$-Tambara functors. Roughly speaking, these may be thought of as generalizations of Galois field extensions $K\rightarrow L$ with Galois group $G$. In this situation, for each subgroup $H$ of $G$ we have a fixed-point intermediate field $L^{H}$, inclusions and conjugations between the various fixed-point fields, residual Weyl group $W_{G}H$ actions on fixed-points $L^{H}$, and Galois-theoretic norm and trace maps between the fixed-point subfields. A Tambara functor may be thought of as relaxing the condition that the inclusions of subfields are injective, and additionally relaxing the assumption that everything in sight is a field. In particular, a Tambara functor $R$ consists of the data of rings $R(G/H)$ for each subgroup $H$ of $G$, with contravariant functoriality along maps of transitive $G$-sets, and covariantly functorial norms and transfers satisfying many conditions (which we recall in Section 2.3). A Green functor is just a Tambara functor without any norms.
The results of the present article arose out of attempts to generalize some of the results of the author’s previous work [Wis24]. Namely, while Tambara functors are equivariant algebra analogues of rings, the notion of a field-like Tambara functor [Nak12] is an equivariant analogue of a field. There is a certain functor, coinduction, denoted by $\mathrm{Coind}_{H}^{G}$, from the category of $H$-Tambara functors to the category of $G$-Tambara functors, and in [Wis24], the author shows that when $G=C_{p^{n}}$ is cyclic of prime power order, every field-like $G$-Tambara functor is the coinduction of a particularly nice kind of field-like $G$-Tambara functor. Coinduction admits a concrete description, is a right adjoint, and enjoys many other nice properties, so this gives a good handle on field-like $C_{p^{n}}$-Tambara functors.
@@
From this we immediately obtain a generalization of one of the main results of [Wis24], from $G=C_{p^{n}}$ a cyclic group of prime power order to arbitrary finite $G$.
@@
This statement is a slightly weaker form of what we actually prove, Theorem 3.11. The norms present in Tambara functors are actually essential in the proof: we observe that Theorem 3.11 fails for Green functors, which are essentially Tambara functors without norms, in Example 4.17. In fact, the failure is pretty catastrophic, as there are multiple distinct ingredients in the proof which make very heavy use of the norms present in a Tambara functor.
@@
We note that coinduced Tambara functors arise “in nature” fairly often, so that the above proposition applies to concrete examples of interest. For example, the main result of [SSW25] roughly says that the algebraically closed fields in the category of $G$-Tambara functors are all coinduced. Additionally, as is well-known to experts, there are many free modules in equivariant algebra which are obtained by forgetting some of the structure of a coinduced Tambara functor (cf. Proposition 5.5).
@@
We warn the reader that induction and coinduction are the same for Mackey functors, but not for Green or Tambara functors. We will not require induction of Green or Tambara functors, so we opt to refer only to coinduction everywhere for uniformity, but whenever we mention the coinduction of a Mackey functor $M$, this is of course the same thing as the induction of $M$.
@@
With this heuristic, we set out to study the interactions between Theorem 3.11 and many different concepts in equivariant algebra. We start with two shorter applications. First, the author intends to investigate affine étale group schemes over Tambara functors in future work, and we give some sample applications along these lines. Second, in Section 5.2 we sketch the topological picture which should be a consequence of Theorem 3.11, under the assumption that the work of Yang [Yan25] generalizes from $G=C_{p}$ to all finite groups $G$. No knowledge of equivariant homotopy theory is required, except for this short subsection on $G$-$\mathbb{E}_{\infty}$ rings.
Next we move on to lengthier applications. Tambara functors are equipped with the notion of a prime ideal, which are organized into the Nakaoka spectrum, an analogue of the Zariski spectrum [Nak12]. We observe that the ideals of $k$ are in bijection with the ideals of $\mathrm{Coind}_{H}^{G}k$, and this bijection induces a homeomorphism of Nakaoka spectra. This allows us to construct a stratification of the Nakaoka spectrum of any $G$-Tambara functor $k$ by subspaces which are the images under certain continuous maps of the Nakaoka spectra of $\mathrm{Res}_{H}^{G}k$. We expect this stratification to be useful for the computation of Nakaoka spectra.
Recall that Nakaoka defines a domain-like Tambara functor as a $G$-Tambara functor $k$ for which the zero ideal is prime.
Theorem H (cf. Proposition 5.27).
A Tambara functor $k$ is domain-like if and only if it is coinduced from a clarified domain-like Tambara functor. We allow the trivial coinduction $\mathrm{Coind}_{G}^{G}$, which is the identity functor, in which case $k$ is itself clarified.
We take the perspective that this result actually motivates the study of domain-like Tambara functors. To the author’s knowledge, not much is known about domain-like Tambara functors besides the fact that the Burnside Tambara functor $\mathcal{A}_{G}$ is domain-like, and the sufficient conditions for a $G$-Tambara functor to be domain-like of Nakaoka [Nak12], and of [CMQ+24] when $G=C_{p}$. On the other hand, our result is actually a property enjoyed by all domain-like Tambara functors. Also, note that this is an additional case of our product decomposition theorem, Theorem 3.11, holding, without the Noetherian assumption.
Theorem I (cf. Proposition 5.17).
There is a bijection between the Tambara ideals of an $H$-Tambara functor $k$ and the Tambara ideals of the $G$-Tambara functor $\mathrm{Coind}_{H}^{G}k$. This bijection respects containment, ideal products, and primality. Thus it determines a homeomorphism of Nakaoka spectra.
To the author’s knowledge, we therefore obtain the first computations of the Nakaoka spectra of $G$-Tambara functors which are not fixed-point and for which $G$ is not abelian. For example, consider the Burnside $C_{p}$-Tambara functor $\mathcal{A}_{C_{p}}$, whose Nakaoka spectrum was computed by Nakaoka [Nak14] (see also [CG23] and [CMQ+24]). We can coinduce along the inclusion into the symmetric group $\Sigma_{p}$, which induces a homeomorphism between the spectrum of $\mathcal{A}_{C_{p}}$ and the spectrum of $\mathrm{Coind}_{C_{p}}^{\Sigma_{p}}\mathcal{A}_{C_{p}}$, which is not a fixed-point Tambara functor. More generally, this reduces the computation of the Nakaoka spectrum of any Tambara functor satisfying the conclusion of Theorem 3.11 to the computation of the Nakaoka spectrum of clarified Tambara functors.
Corollary J (cf. Theorem 5.19).
Let $k$ be a $G$-Tambara functor such that either $k(G/e)$ is Noetherian or $k$ is domain-like. Then the Nakaoka spectrum of $k$ is homeomorphic to a disjoint union of Nakaoka spectra of clarified Tambara functors.
For another application of the clarification picture when $G=C_{p}$, we observe that the our results interact nicely with the story of prisms and transversal $C_{p}$-Tambara functors developed by Sulyma [Sul23].
Finally, we conclude with our lengthiest application, the study of Nullstellensatzian objects in the category of clarified Tambara functors. Introduced in [BSY22], Nullstellensatzian objects are generalizations of algebraically closed fields, by abstraction of Hilbert’s Nullstellensatz. Previous work of the author joint with Schuchardt and Spitz classifies the Nullstellensatzian $G$-Tambara functors: they are precisely the coinductions $\mathrm{Coind}_{e}^{G}\mathbb{F}$ of algebraically closed fields. In fact, we are able to give a short proof of this result with the machinery we build up, which uses the same ideas as one of the two proofs given in [SSW25].
@@
In light of this, there are two natural questions. First, 5.48 asserts that there are field-like Nullstellensatzian clarified $G$-Tambara functors of every form not eliminated by Theorem 5.45 and Artin-Schreier theory. Second, 5.47 asks what the remaining Nullstellensatzian clarified $G$-Tambara functors look like. We are able to make slight progress along this second front.
1.1. Contents
The contents of this article are organized as follows. We begin in Section 2 with a review of the necessary background on Mackey, Green, and Tambara functors for the main results of the next two sections.
In Section 3 we extract many useful consequences of the Tambara functor axioms, and prove our main theorem.
Section 4 concerns an analysis of morphisms between Tambara functors admitting a product decomposition as in Theorem 3.11. Additionally, we elaborate on some of the unusual behavior of coproducts in the category of clarified Tambara functors.
Finally, Section 5 contains all of our applications. In order, we discuss affine group schemes over Tambara functors, a topological version of the results of Section 3, Morita invariance, Nakaoka spectra, domain-like Tambara functors, prisms and transversal Tambara functors, and finally Nullstellensatzian clarified Tambara functors.
1.2. Acknowledgments
The author thanks Mike Hill for a wealth of insightful suggestions and David Chan for many helpful conversations. Additionally, the author thanks David Mehrle, Natalie Stewart, Maxime Ramzi, and the wonderful participants of Northwestern’s informal Number Theory seminar organized by Haochen Cheng for insightful suggestions. For providing comments and suggestions on early drafts, the author thanks David Mehrle and Jackson Morris. Finally, the author thanks Alley Koenig for suggesting the term “clarified” as a replacement for “pure”.
@@ 2.1. Mackey functors
[…]
with composition induced by the pullback of finite $G$-sets.
Definition 2.2.
A semi-Mackey functor is a product-preserving functor $\mathrm{Span}(G-\mathsf{Set}^{\mathrm{fin}})\rightarrow\mathsf{Set}$ from the Burnside category to sets. A morphism of semi-Mackey functors is a natural transformation.
Given a finite $G$-set $X$ and semi-Mackey functor $M$, we refer to $M(X)$ as the value of $M$ in level $X$. If $X\cong G/H$, we may refer to $M(X)$ as the value of $M$ in level $H$. Given a map of $G$-sets $f:X\rightarrow Y$, there are two natural bispans one may form: $X\leftarrow X\rightarrow Y$ and $X\leftarrow Y\rightarrow Y$. The image of the former morphism under $M$ is called transfer and the image of the latter under $M$ is called restriction.
Products and coproducts in the Burnside category are given by disjoint union. From this, one may deduce that semi-Mackey functors naturally land in the category of commutative monoids. In fact, the addition in the commutative monoid structure is induced by transfer along the fold map $X\sqcup X\rightarrow X$.
Definition 2.3.
A Mackey functor is a semi-Mackey functor $M$ for which each commutative monoid $M(X)$ is an abelian group. A morphism of Mackey functors is a morphism of underlying semi-Mackey functors. The category of $G$-Mackey functors is denoted by $G$-$\mathsf{Mack}$.
Given a semi-Mackey functor, post-composition with group completion therefore yields a Mackey functor.
@@
The following alternative description of a Mackey functor is well-known.
Proposition 2.6.
A Mackey functor $M$ is precisely the following data.
  1. (1)
    For each subgroup $H\subset G$, an abelian group $M(G/H)$.
  2. (2)
    For each $g\in G$, an isomorphism $c_{g}:M(G/H)\rightarrow M(G/gHg^{-1})$.
  3. (3)
    For each subgroup inclusion $K\subset H$, a restriction morphism
    $\mathrm{Res}_{K}^{H}:M(G/H)\rightarrow M(G/K)$
    such that $\mathrm{Res}_{gKg^{-1}}^{H}c_{g}=c_{g}\mathrm{Res}_{K}^{H}$ and $\mathrm{Res}_{K}^{H}\mathrm{Res}_{L}^{K}=\mathrm{Res}_{L}^{H}$.
  4. (4)
    For each subgroup inclusion $K\subset H$, a transfer morphism
    $\mathrm{Tr}_{K}^{H}:M(G/K)\rightarrow M(G/H)$
    such that $\mathrm{Tr}_{gKg^{-1}}^{H}c_{g}=c_{g}\mathrm{Tr}_{K}^{H}$ and $\mathrm{Tr}_{K}^{H}\mathrm{Tr}_{L}^{K}=\mathrm{Tr}_{L}^{H}$.
  5. (5)
    Restrictions and transfers satisfy the double coset formula:
    $\mathrm{Res}_{L}^{H}\mathrm{Tr}_{K}^{H}=\sum_{g\in L\backslash H/K}\mathrm{Tr}_{L\cap gKg^{-1}}^{L}c_{g}\mathrm{Res}_{gLg^{-1}\cap K}^{K}$
A similar description holds for semi-Mackey functors. Given a Mackey functor or semi-Mackey functor, we sometimes regard it as only taking values on transitive $G$-sets, and sometimes regard it as taking values on all finite $G$-sets. We can convert between these two perspectives by either restricting to transitive $G$-sets or extending in the unique way which preserves products.
With this description, a morphism of Mackey functors $M\rightarrow N$ is given by a morphisms $M(G/H)\rightarrow N(G/H)$ for each $H\subset G$ which intertwine conjugation $c_{g}$, restriction $\mathrm{Res}_{K}^{H}$, and transfer $\mathrm{Tr}_{K}^{H}$.
Corollary 2.7.
If all restrictions in a Mackey functor are injective, then the transfers are totally determined by the double coset formula.
The isomorphisms $c_{g}$ induce an action of the Weyl group $W_{G}H:=N_{G}H/H$ on $M(G/H)$.
Example 2.8.
The assignment $G/H\mapsto\mathrm{RU}(H)$ sending $G/H$ to the complex representation ring of $H$ defines a $G$-Mackey functor $\mathrm{RU}_{G}$. In fact, we have already seen this example as we have an isomorphism $\mathrm{RU}_{G}\cong\underline{\pi}_{0}(KU_{G})$ determined by
$RU(H)\cong\pi_{0}^{H}(KU_{H})\cong\pi_{0}^{G}(G/H_{+}\wedge KU_{G})$
where $KU_{G}$ denotes $G$-equivariant complex topological $K$-theory.
Example 2.9.
Let $G$ act on an abelian group $D$. The assignment $G/H\mapsto D^{H}$ defines a Mackey functor $\mathrm{FP}(D)$, the fixed point Mackey functor associated to ($G$ acting on) $D$. Restrictions are given by inclusions of fixed points, Weyl actions are given by the residual Weyl action on fixed points, and the transfers are totally determined by the double coset formula.
Definition 2.10.
The box product $\boxtimes$ of two Mackey functors is given by the Day convolution product with respect to disjoint union of finite $G$-sets and tensor product of abelian groups.
The box product makes the category of Mackey functors into a symmetric monoidal category with unit the Burnside Mackey functor. The box product enjoys a universal property given by Dress pairings which are analogous to the universal property enjoyed by the tensor product.
Proposition 2.11.
Let $M$, $N$, and $K$ be Mackey functors. A morphism
$M\boxtimes N\rightarrow K$
is precisely the data of, for each $H\subset G$, an abelian group homomorphism $f_{H}:M(G/H)\otimes N(G/H)\rightarrow K(G/H)$, satisfying
  1. (1)
    $\mathrm{Res}_{H}^{L}\circ f_{H}=f_{L}\circ\left(\mathrm{Res}_{H}^{L}\otimes\mathrm{Res}_{H}^{L}\right)$,
  2. (2)
    $\mathrm{Tr}_{H}^{L}\circ f_{H}\circ\left(Id\otimes\mathrm{Res}_{H}^{L}\right)=f_{L}\circ\left(\mathrm{Tr}_{H}^{L}\otimes Id\right)$, and
  3. (3)
    $\mathrm{Tr}_{H}^{L}\circ f_{H}\circ\left(\mathrm{Res}_{H}^{L}\otimes Id\right)=f_{L}\circ\left(Id\otimes\mathrm{Tr}_{H}^{L}\right)$.
The value of the box product at the bottom level, $G/e$, is particularly pleasant. Namely, we have $(M\boxtimes N)(G/e)\cong M(G/e)\otimes N(G/e)$, with the Weyl group $G\cong W_{G}e$ acting diagonally.
We recall a few more standard facts about Mackey functors.
Proposition 2.12.
The category of Mackey functors is bicomplete. Limits are computed levelwise (that is, the functors $M\mapsto M(G/H)$ preserves limits) and filtered colimits are computed levelwise.
Proposition 2.13.
The category of Mackey functors is an abelian category. Kernels and cokernels are computed levelwise.
There are some useful functors which relate the categories of $G$-Mackey functors for varying $G$, like restriction, induction, and coinduction. We review these in a later subsection.
2.2. Green functors
@@
A Green functor $R$ is a Mackey functor which is a commutative monoid with respect to the box product. A morphism of Green functors is a morphism which respects the monoid structure. The category of $G$-Green functors is denoted by $G$-$\mathsf{Green}$.
@@
Example 2.16.
Let $E$ be a homotopy commutative ring (genuine) $G$-spectrum. Then the Mackey functor $\underline{\pi}_{0}(E)$ inherits the structure of a Green functor from the multiplication on $E$.
There is a concrete description of a Green functor obtained by unraveling the Dress pairing description of the box product.
Proposition 2.17.
A Green functor is the same data as a Mackey functor $R$ and multiplication maps for each $H\subset G$ making $R(G/H)$ into a (unital commutative) ring, subject to the following conditions:
  1. (1)
    the restriction and conjugation maps are ring homomorphisms,
  2. (2)
    the transfer and restriction maps satisfy the Frobenius reciprocity relations
    $\mathrm{Tr}^{K}_{H}(x)\cdot y=\mathrm{Tr}^{K}_{H}(x\cdot\mathrm{Res}^{K}_{H}(y))\quad\mathrm{and}\quad x\cdot\mathrm{Tr}^{K}_{H}(y)=\mathrm{Tr}^{K}_{H}(\mathrm{Res}^{K}_{H}(x)\cdot y)$
    whenever these make sense.
As with Mackey functors, we sometimes regard a Green functor as indexed only on transitive $G$-sets, and sometimes as indexed on all finite $G$-sets by taking disjoint unions of transitive $G$-sets to the appropriate products of rings.
Example 2.18.
Let $G$ act on a ring $D$. Then the fixed-point Mackey functor $\mathrm{FP}(D)$ inherits the structure of a Green functor from the ring structure on $D$.
Example 2.19.
The Mackey functor $\mathrm{RU}_{G}$ is a Green functor via the ring structures on $\mathrm{RU}(H)$ induced by the tensor product of $H$-representations. Equivalently, it is a Green functor via the isomorphism $\mathrm{RU}_{G}\cong\underline{\pi}_{0}(KU_{G})$ and the standard ring structure on $KU_{G}$.
A morphism of Green functors $R\rightarrow S$ is therefore a morphism of underlying Mackey functors which is given levelwise by ring maps $R(G/H)\rightarrow S(G/H)$.
Additionally, one may define modules over a Green functor in terms of the box product. The Dress pairing description of the box product again allows one to give a concrete definition.
@@
Proposition 2.21.
A module over a Green functor $R$ consists of a Mackey functor $M$ with action maps $R(G/H)\otimes M(G/H)\rightarrow M(G/H)$ expressing $M(G/H)$ as an $R(G/H)$-module subject to the following conditions:
  1. (1)
    for any $H\leq K$ we have $\mathrm{Res}^{K}_{H}(rm)=\mathrm{Res}^{K}_{H}(r)\cdot\mathrm{Res}^{K}_{H}(m)$,
  2. (2)
    for any $H\leq G$ and $g\in G$ we have $c_{g}(rm)=c_{g}(r)c_{g}(m)$,
  3. (3)
    the transfer and restriction maps satisfy the Frobenius reciprocity relations
    $\mathrm{Tr}^{K}_{H}(r)\cdot m=\mathrm{Tr}^{K}_{H}(r\cdot\mathrm{Res}^{K}_{H}(m))\quad\mathrm{and}\quad r\cdot\mathrm{Tr}^{K}_{H}(m)=\mathrm{Tr}^{K}_{H}(\mathrm{Res}^{K}_{H}(r)\cdot m)$
    whenever these make sense.
A morphism $M\rightarrow N$ of $R$ modules therefore consists of a morphism of Mackey functors which is given levelwise by a morphism $M(G/H)\rightarrow N(G/H)$ of $R(G/H)$-modules.
@@
Proposition 2.23.
Let $M$, $N$, and $K$ be modules over a Green functor $R$. A morphism
$M\boxtimes_{R}N\rightarrow K$
is precisely the data of, for each $H\subset G$, an $R(G/H)$-module homomorphism
$f_{H}:M(G/H)\otimes_{R(G/H)}N(G/H)\rightarrow K(G/H)$
satisfying the three Dress pairing conditions of Proposition 2.11.
@@
2.3. Tambara functors
It turns out that Green functors do not encode all of the multiplicative structure naturally adherent to $\pi_{0}$ of most ring genuine $G$-spectra which occur in nature. In fact, the same kind of extra structure is present also on fixed-point Green functors and the Burnside Mackey functor. This extra structure is encoded by norms, which are like multiplicative analogues of transfers. A Green functor with norms is called a Tambara functor, as introduced by Tambara [Tam93] under the name “$TNR$ functors”. It turns out that Tambara functors are described in terms of bispans, analogous to the description of Mackey functors in terms of spans.
Definition 2.26.
The category $\mathrm{bispan}(G-\mathsf{Set}^{\mathrm{fin}})$ of bispans of finite $G$-sets has objects the category of finite $G$-sets and morphisms $X\rightarrow Y$ equivalence classes of diagrams $\phi:X\leftarrow A\rightarrow B\rightarrow Y$ under the relation $\phi\sim\phi^{\prime}$ if there is a commutative diagram
[…]
with composition defined as in [Tam93].
As before, coproducts and products in $\mathrm{bispan}(G-\mathsf{Set}^{\mathrm{fin}})$ are given by disjoint union of $G$-sets. Any product-preserving functor out of $\mathrm{bispan}(G-\mathsf{Set}^{\mathrm{fin}})$ therefore naturally takes values in commutative monoids in the target category (with respect to the symmetric monoidal structure given by the product).

### v2

@@ Abstract.
Tambara functors are equivariant analogues of rings arising in representation theory and equivariant homotopy theory. We introduce the notion of a clarified Tambara functor and show that under mild conditions every Tambara functor admits a decomposition as a product of coinductions of clarified Tambara functors; projection onto the non-coinduced part defines a reflective localization we call clarification. Through this perspective we study Morita invariance and $K$-theory of Tambara functors, field-like Tambara functors, and Nullstellensatzian clarified Tambara functors.
@@ 1. Introduction
We focus our attention on $G$-Tambara functors. Roughly speaking, these may be thought of as generalizations of Galois field extensions $K\rightarrow L$ with Galois group $G$. In this situation, for each subgroup $H$ of $G$ we have a fixed-point intermediate field $L^{H}$, inclusions and conjugations between the various fixed-point fields, residual Weyl group $W_{G}H$ actions on fixed-points $L^{H}$, and Galois-theoretic norm and trace maps between the fixed-point subfields. A Tambara functor may be thought of as relaxing the condition that the inclusions of subfields are injective, and additionally relaxing the assumption that everything in sight is a field. In particular, a Tambara functor $R$ consists of the data of rings $R(G/H)$ for each subgroup $H$ of $G$, with contravariant functoriality along maps of transitive $G$-sets, and covariantly functorial norms and transfers satisfying many conditions. A Green functor is just a Tambara functor without any norms.
The results of the present article arose out of attempts to generalize some of the results of the author’s previous work [Wis25b]. Namely, while Tambara functors are equivariant algebra analogues of rings, the notion of a field-like Tambara functor [Nak12] is an equivariant analogue of a field. There is a certain functor, coinduction, denoted by $\mathrm{Coind}_{H}^{G}$, from the category of $H$-Tambara functors to the category of $G$-Tambara functors, and in [Wis25b], the author shows that when $G=C_{p^{n}}$ is cyclic of prime power order, every field-like $G$-Tambara functor is the coinduction of a particularly nice kind of field-like $G$-Tambara functor. Coinduction admits a concrete description, is a right adjoint, and enjoys many other nice properties, so this gives a good handle on field-like $C_{p^{n}}$-Tambara functors.
@@
From this we immediately obtain a generalization of one of the main results of [Wis25b], from $G=C_{p^{n}}$ a cyclic group of prime power order to arbitrary finite $G$.
@@
The norms present in Tambara functors are actually essential in the proof: we observe that Theorem 3.10 fails for Green functors, which are essentially Tambara functors without norms, in Example 4.17. In fact, the failure is pretty catastrophic, as there are multiple distinct ingredients in the proof which make very heavy use of the norms present in a Tambara functor.
@@
We note that coinduced Tambara functors arise “in nature” fairly often, so that the above proposition applies to concrete examples of interest. For example, the main result of [SSW25] roughly says that the algebraically closed fields in the category of $G$-Tambara functors are all coinduced. Additionally, in [Wis25a, Theorem A] it is shown that the adjunction units $k\rightarrow\mathrm{CoInd}_{H}^{G}\mathrm{Res}_{H}^{G}k$ are all étale (and in some cases these are essentially all of the finite étale $k$-algebras, cf. [Wis25a, Theorem C]).
@@
We warn the reader that induction and coinduction are the same for Mackey functors, but not for Green or Tambara functors. We will not require induction of Green or Tambara functors, so we opt to refer only to coinduction everywhere for uniformity, but whenever we mention the coinduction of a Mackey functor $M$, this is the same thing as the induction of $M$.
@@
These are evidently fairly useful results: in [Wis25a, Theorem C], the author uses Theorem 3.10 to classify finite étale extensions of the constant $G$-Tambara functor $\underline{\mathbb{F}}$ associated to any algebraically closed field $\mathbb{F}$, which leads to a complete description of affine étale group schemes over $\underline{\mathbb{F}}$ in [Wis25a, Corollary D]. On the other hand, in [Wis25c], the author uses Corollary 5.9 and Theorem 5.6 to show that the Nakaoka spectrum of $\mathrm{CoInd}_{H}^{G}R$ is naturally homeomorphic to the Nakaoka spectrum of $R$ for $R$ any $H$-Tambara functor and to construct a stratification of the Nakaoka spectrum of any Tambara functor (cf. [Wis25c, Theorem A, Definition B]).
Finally, we conclude with a study of Nullstellensatzian objects in the category of clarified Tambara functors. Introduced in [BSY22], Nullstellensatzian objects are generalizations of algebraically closed fields, by abstraction of Hilbert’s Nullstellensatz. Previous work of the author joint with Schuchardt and Spitz classifies the Nullstellensatzian $G$-Tambara functors: they are precisely the coinductions $\mathrm{Coind}_{e}^{G}\mathbb{F}$ of algebraically closed fields. In fact, we are able to give a short proof of this result with the machinery we build up, which uses the same ideas as one of the two proofs given in [SSW25].
@@
In light of this, there are two natural questions. First, 5.21 asserts that there are field-like Nullstellensatzian clarified $G$-Tambara functors of every form not eliminated by Theorem 5.18 and Artin-Schreier theory. Second, 5.20 asks what the remaining Nullstellensatzian clarified $G$-Tambara functors look like. We are able to make slight progress along this second front.
Contents. The contents of this article are organized as follows. We begin in Section 2 with a review of the necessary background on Mackey, Green, and Tambara functors for the main results of the next two sections. In Section 3 we extract many useful consequences of the Tambara functor axioms, and prove our main theorem.
Section 4 concerns an analysis of morphisms between Tambara functors admitting a product decomposition as in Theorem 3.10. Additionally, we elaborate on some of the unusual behavior of coproducts in the category of clarified Tambara functors. Finally, Section 5 contains all of our applications. Beginning with an overview of applications appearing in other article, Section 5.1 studies modules and algebras over coinductions, and Section 5.2 studies Nullstellensatzian clarified Tambara functors.
Acknowledgments. The author thanks Mike Hill for a wealth of insightful suggestions, including a major rearranging of the content of the early versions of this article, and David Chan for many helpful conversations. Additionally, the author thanks David Mehrle, Natalie Stewart, Maxime Ramzi, and the wonderful participants of Northwestern’s informal Number Theory seminar organized by Haochen Cheng for helpful suggestions. For providing comments and suggestions on early drafts, the author thanks Mike Hill, David Mehrle, Jackson Morris, Birgit Richter, and Yuri Sulyma. Finally, the author thanks Alley Koenig for suggesting the term “clarified” as a replacement for “pure”.
@@ 2. Review of equivariant algebra
A Mackey functor is a product-preserving functor $M:\mathrm{Span}(G-\mathrm{Set}^{\mathrm{fin}})\rightarrow\mathrm{Set}$ from the Burnside category to sets, for which each monoid $M(X)$ is actually a group. A morphism of Mackey functors is a natural transformation.
Given a finite $G$-set $X$ and semi-Mackey functor $M$, we refer to $M(X)$ as the value of $M$ in level $X$. Given a map of $G$-sets $f:X\rightarrow Y$, there are two natural bispans one may form: $X\leftarrow X\rightarrow Y$ and $Y\leftarrow X\rightarrow X$. The image of the former morphism under $M$ is called transfer and the image of the latter under $M$ is called restriction.
@@
The box product $\boxtimes$ of two Mackey functors is given by the Day convolution product with respect to the product of finite $G$-sets and tensor product of abelian groups. The box product makes the category of Mackey functors into a symmetric monoidal category with unit the Burnside Mackey functor. The value of the box product at the bottom level, $G/e$, is particularly pleasant. Namely, we have $(M\boxtimes N)(G/e)\cong M(G/e)\otimes N(G/e)$, with the Weyl group $G\cong W_{G}e$ acting diagonally.
@@
A Green functor $R$ is a Mackey functor which is a commutative monoid with respect to the box product. A morphism of Green functors is a morphism of Mackey functors which respects the monoid structure. The category of $G$-Green functors is denoted by $G$-$\mathrm{Green}$.
@@
Additionally, one may define modules over a Green functor in terms of the box product.
@@
Tambara functors are Green functors with norms, multiplicative analogues of transfers, and were first introduced by Tambara in [Tam93] under the name $TNR$-functors.